OpenStudy (sleepyjess):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 3, -13, and 5 + 4i

OpenStudy (sleepyjess):

@dumbcow

OpenStudy (anonymous):

3 roots but one is a complex root. complex roots come in conjugate pairs if the coefficients of the polynomial are real. so that means you need a 4th degree polynomial with real coefficients to satisfy the requirements.

OpenStudy (sleepyjess):

Do you want to see the answer choices or does it matter?

OpenStudy (anonymous):

so your roots need to be 3, -13, 5 + 4i and 5 - 4i. that menas your factors should be... (x-3)(x+13)(x- 5 - 4i)(x - 5+4i)

OpenStudy (anonymous):

multiply those out to get your answer.

OpenStudy (sleepyjess):

f(x) = x^4 - 98x^2 + 800x - 1599

OpenStudy (dumbcow):

for imaginary roots, i like to use this method: x = 5 + 4i x-5 = 4i (x-5)^2 = (4i)^2 x^2 -10x +25 = -16 x^2 -10x +41 = 0 polynomial: ---> (x-3)(x+13)(x^2 -10x+41)

OpenStudy (sleepyjess):

Could you help me on 2 more @dumbcow ?

OpenStudy (dumbcow):

ok